# Approaching Rate-Distortion Limits in Neural Compression with Lattice Transform Coding

### v1

@@ 2 Related Work
which is a scale-invariant measure of the lattice’s packing efficiency. Recent work describing lattices which have the lowest known NSM can be found in (Agrell and Allen, 2023; Lyu et al., 2022; Agrell et al., 2024).
Another challenge is the quantization (also referred to as decoding) operation itself $Q_{\Lambda}(\bm{x})$, which requires solving the closest-vector problem (CVP); this is NP-hard in general. For many of the best-known lattices in dimensions up to 24 ($n=2$: hexagonal; $n=4$: $D_{n}^{*}$,; $n=8$: Gosset $E_{8}$; $n=16$: $\Lambda_{16}$ Barnes-Wall; and $n=24$: $\Lambda_{24}$ Leech) efficient algorithms to solve CVP are known (Conway and Sloane, 1999). For general lattices, CVP can be solved using the Schnorr-Euchner (SE) decoder (Agrell et al., 2002), but solving exact CVP using SE is much slower compared to custom lattice decoders, and is not as easily parallelizable across batches for GPU implementation.
In this paper, we use the CVP algorithms outlined in Conway and Sloane (1999, Ch. 20). For the Barnes-Wall lattice $\Lambda_{16}$ in dimension 16, we use the algorithm outlined in Conway and Sloane (1984). Finally, the Leech lattice $\Lambda_{24}$ in dimension 24 has the slower method described in Conway and Sloane (1984) and the improved version in Conway and Sloane (1986).
Companding.
Companding has typically been analyzed in the (asymptotically) large rate, small distortion regime. Optimal scalar quantization can be implemented using scalar companding (Bennett, 1948). For the vector case, vector companders cannot implement optimal VQ (Gersho, 1979; Bucklew, 1981; 1983), except for a very small class of sources. Despite this, it is still possible optimal companding may perform near-optimal VQ.
This is shown in Linder et al. (1999), who show that the optimal compander can arbitrarily approach the rate-distortion function in the limit of large dimensions and rates:
$\lim_{D\rightarrow 0}\frac{1}{n}\left(H(Q_{D,f})-R_{\bm{x}}(D)\right)$ $=\frac{1}{2}\log(2\pi eL_{\Lambda}),$ (5)
where $R_{\bm{x}}(D)$ is the rate-distortion function of $\bm{x}\sim P_{\bm{x}}$, $H(Q_{D,f})$ is the entropy of the optimal compander at distortion $D$, and $L_{\Lambda}$ is the NSM for a lattice $\Lambda$ in $n$ dimensions. As the authors note, if $G_{n}$ is the minimal NSM of any lattice in $n$ dimensions, $\lim_{n\rightarrow\infty}G_{n}=\frac{1}{2\pi e}$, such that (5) converges to 0 at a rate $O\left(\frac{\log n}{n}\right)$. This result indicates that the gap to $R_{\bm{x}}(D)$ is determined by the packing efficiency of the lattice and not $f$; for the integer lattice, the NSM is $1/12$, which yields a gap of $\approx 0.255$ bits per dimension. This indicates that it is impossible for NTC (which uses the integer lattice) to be optimal with respect to the rate-distortion function.
The companding literature does not address non-asymptotic rate regimes, which NTC models typically operate at. In this paper, we provide experimental evidence that the gains of vector companding still hold for non-asymptotic rates.
3 NTC is Sub-Optimal on i.i.d. Sequences
Let $P_{S}\in\mathcal{P}(\mathbb{R})$ be a one-dimensional distribution, and suppose that the entries of $\bm{x}$ are i.i.d. samples $S_{1},\dots,S_{n}\stackrel{{\scriptstyle\textrm{i.i.d.}}}{{\sim}}P_{S}$. We first choose $P_{S}=\mathcal{N}(0,1)$, so that $\bm{x}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$. This is because i.i.d. Gaussians are the most difficult source to compress, and have no low-dimensional latent structure for the transforms. We train an NTC model following Ballé et al. (2021) using a factorized entropy model for the distribution of the latent $\bm{y}$. We use the objective in (1), with $\operatorname{\mathsf{d}}(\bm{x},\hat{\bm{x}})=\|\bm{x}-\hat{\bm{x}}\|_{2}^{2}$, and normalize the entropy and distortion terms by $n$ to compare with the rate-distortion function of $S$, as well as entropy-constrained vector quantization (ECVQ) which is the optimal quantizer for any fixed $n$.
Shown in Fig. 1, NTC with $n=1$ is able to recover ECVQ with $n=1$, also known as entropy-constrained scalar quantization (ECSQ). This is perhaps expected, since the task of the transforms is to scale the 1-D Gaussian to be optimally quantized on an integer grid in 1-D, and then scale the quantized latent to form the reconstruction. As expected, a gap between this (optimal) one-shot performance of coding $S$ and $R(D)$ of $S$ exists.
For $n=2$, however, NTC is unable to recover ECVQ, and always achieves the same performance as that of ECSQ. To see this more clearly, we define the normalized rate-redundancy of a compressor $\mathsf{c}$ as
$\mathcal{E}_{\mathsf{c}}(D):=\frac{R_{\mathsf{c}}(D)-R(D)}{R(D)},$ (6)
where $R_{\mathsf{c}}(D)$ is the rate achieved by $\mathsf{c}$ at distortion $D$, and $R(D)$ is the rate-distortion function of $S$. Intuitively, $\mathcal{E}_{\mathsf{c}}(D)$ measures the fraction of additional bits per sample required to code a sequence relative to $R(D)$.
Figure 3: $\mathcal{E}_{\mathsf{c}}(D)$ of NTC for $n=1,2$.
(a) NTC.
(b) ECVQ.
Figure 4: Quantization regions of i.i.d. Gaussians with $n=2$.
Shown in Fig. 3, we see that the $\mathcal{E}_{\mathsf{c}}(D)$ for ECVQ at $n=2$ is lower than that of ECSQ. However, NTC at $n=2$ is unable to outperform $\mathcal{E}_{\mathsf{c}}(D)$ of ECSQ, which is the lower bound of $\mathcal{E}_{\mathsf{c}}(D)$ at $n=1$. Visualizing the quantization regions and centers on the 2-D Gaussian distribution (Fig. 3(a)) brings more clarity as to why this is the case. It is clear that NTC is merely implementing a scalar quantization of the two dimensional input, perhaps with a rotation, which is equivalent to applying ECSQ to the entries of $\bm{x}$ independently. If we compare to the optimal ECVQ in $n=2$, shown in Fig. 3(b), we see that the optimal quantization regions are hexagonal, not square.
In our experiments, we varied the depth, width, nonlinearities, biases, and even increased the latent dimension to a higher-dimensional space, all to no avail. Setting $n>2$ did not help either. All models recovered the performance of ECSQ at best. Thus, it appears that $g_{a}$ and $g_{s}$ parameterized as neural networks are unable to generate optimal quantization regions in the original space when scalar quantization is used in the latent space. This experimentally verifies results in companding (Linder et al., 1999).
4 Lattice Transform Coding
We seek a compressor that can provide better performance than scalar quantization yet avoid the high complexity of vector quantization. A lattice quantizer $Q_{\Lambda}$ provides an intermediate solution that manages both complexity and space-filling efficiency. We first discuss how to optimize the rate-distortion loss under lattice quantization in the latent space, which brings several challenges. Then we the role of the entropy model in LTC. Lastly, we discuss Block-LTC, which uses LTC to implement a block coding scheme for i.i.d. vector sources.
4.1 Rate-Distortion Loss Under Lattice Quantization
Typically, the objective in (1) is replaced with a proxy loss since quantization operations are not differentiable. We investigate both the straight-through estimator (STE) and dithered quantization with uniform noisy proxy. The STE approach can use (1) directly as an objective during training, but may lead to training instabilities. The uniform noise proxy has stabler training, but requires a rate-distortion proxy loss of (1) using a connection to dithered quantization.
We assume that the lattice quantizer $Q_{\Lambda}$ under consideration has an efficient algorithm to find the closest vector; see the appendix for lattice details. We also assume that the lattice’s generator matrix $\bm{G}$ has been normalized to be unit volume, i.e. $\mathrm{Vol}(V(\bm{0}))=\sqrt{\det(\bm{G}\bm{G}^{\top})}=1$. If a non-unit volume generator matrix $\bm{G}^{\prime}$ is given, one can simply make it unit volume via $\bm{G}=(\det(\bm{G}^{\prime}\bm{G}^{\prime\top}))^{-\frac{1}{2k}}\bm{G}^{\prime}$, where $k$ is the number of rows of $\bm{G}^{\prime}$.
Distortion term.
We use straight-through estimation (STE) for the distortion term $\operatorname{\mathbb{E}}_{\bm{x}}[\operatorname{\mathsf{d}}(\bm{x},g_{s}(Q_{\Lambda}(g_{a}(\bm{x}))))]$. Specifically, we compute the quantized output $\hat{\bm{y}}=Q_{\Lambda}(g_{a}(\bm{x}))$ during the forward pass via
Rate term.
For the rate term, we typically have some parameterized density model $p_{\bm{y}}$ for the continuous latent $\bm{y}$ (more details on this in the next section). This induces a PMF over the quantized latents $p_{\hat{\bm{y}}}$, given by
@@ 4.1 Rate-Distortion Loss Under Lattice Quantization
where $V(\hat{\bm{y}}):=\{\bm{y}:Q_{\Lambda}(\bm{y})=\hat{\bm{y}}\}$. In NTC, where $Q_{\Lambda}$ implements the integer lattice, (8) is easy to integrate when $p_{\bm{y}}$ is assumed factorized (Ballé et al., 2018). In that case, (8) can be written in terms of the CDF of $p_{\bm{y}}$, which is one reason why the factorized $p_{\bm{y}}$ is typically parameterized by its CDF in NTC. However, for lattice quantization, solving the integral in (8) is not as straightforward.
As a result, we compute (8) via Monte-Carlo estimate,
@@
$=\mathrm{Vol}(V(\hat{\bm{y}}))\cdot\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(V(\hat{\bm{y}}))}[p_{\bm{y}}(\bm{u})]$ (10)
$=\mathbb{E}_{\bm{u}^{\prime}\sim\mathrm{Unif}(V(\bm{0}))}[p_{\bm{y}}(\hat{\bm{y}}+\bm{u}^{\prime})],$ (11)
where $p_{\bm{u}}$ is the uniform density over $V(\hat{\bm{y}})$, $\mathrm{Vol}(V(\hat{\bm{y}}))=1$ since we assumed unit-volume lattices, and the last equality holds because for a lattice quantizer $Q_{\Lambda}(\bm{x}+\bm{v})=Q_{\Lambda}(\bm{x})+\bm{v}$, $\forall\bm{v}\in\Lambda$. From Conway and Sloane (1984), we can sample $\bm{u}^{\prime}\sim\mathrm{Unif}(V(\bm{0}))$ by computing $\bm{u}^{\prime}=\tilde{\bm{u}}-Q_{\Lambda}(\tilde{\bm{u}})$, where $\tilde{\bm{u}}=\bm{s}\bm{G}$, and $\bm{s}\sim\mathrm{Unif}([0,1])$. Thus, (11) can be estimated via a sample mean. With the ability to reliably estimate $p_{\hat{\bm{y}}}(\hat{\bm{y}})$, we can also use STE for backpropagation of the rate term $\operatorname{\mathbb{E}}_{\bm{x}}[-\log p_{\hat{\bm{y}}}(\hat{\bm{y}})]$ by using (7).
Dithering.
While STE could be used in both cases, prior works have achieved benefits from using the dithering approach (Ballé et al., 2021). In this case, a random dither $\bm{u}^{\prime}\sim\mathrm{Unif}(V(0))$ is applied prior to quantization, with the resulting rate $\operatorname{\mathbb{E}}_{\bm{x},\bm{u}^{\prime}}[-\log p_{\hat{\bm{y}}}(Q_{\Lambda}(\bm{y}-\bm{u}^{\prime}))]$. This rate is equivalent in the integer lattice to $\operatorname{\mathbb{E}}_{\bm{x},\bm{u}^{\prime}}[-\log p_{\bm{y}+\bm{u}^{\prime}}(\bm{y}+\bm{u}^{\prime})]$ (Ballé et al., 2021). The density $p_{\bm{y}+\bm{u}^{\prime}}$ is a convolved density with likelihoods given by
@@
$=\operatorname{\mathbb{E}}_{\bm{u}^{\prime}\sim\mathrm{Unif}(V(0))}[p_{\bm{y}}(\bm{y}^{\prime}-\bm{u}^{\prime})],$ (13)
which can be estimated via Monte-Carlo.
To apply this for our lattice quantization case during training, rather than provide the STE of the quantized output $\hat{\bm{y}}$ to the Monte-Carlo likelihood estimate, we provide the noise-added continuous latent $\bm{y}+\bm{u}^{\prime}$, where $\bm{u}^{\prime}$ is uniformly distributed over the Voronoi region. In our experiments, we observed that STE was less stable than the dithering approach, but had better performance at low rates, as will be shown in Sec. 5.
4.2 Entropy Models
Previously, the rate (under non-dithered quantization) has been referred to as $\operatorname{\mathbb{E}}_{\bm{x}}[-\log p_{\hat{\bm{y}}}(\hat{\bm{y}})]$, which is actually the cross-entropy between the PMF $P(\hat{\bm{y}}):=\Pr_{\bm{x}}(Q_{\Lambda}(g_{a}(\bm{x})=\hat{\bm{y}})$ on the centers $\hat{\bm{y}}$ induced by $\bm{x}$ and $p_{\hat{\bm{y}}}(\hat{\bm{y}})$, which is some PMF that is learned. Thus, what we call the rate is actually an upper bound on the true entropy of $P(\hat{\bm{y}})$, i.e.
$\operatorname{\mathbb{E}}_{\bm{x}}[-\log P(\hat{\bm{y}})]$ $\leq\operatorname{\mathbb{E}}_{\bm{x}}[-\log P(\hat{\bm{y}})]+D_{\mathsf{KL}}(P(\hat{\bm{y}})||p_{\hat{\bm{y}}}(\hat{\bm{y}}))$
$=\operatorname{\mathbb{E}}_{\bm{x}}[-\log p_{\hat{\bm{y}}}(\hat{\bm{y}})],$ (14)
with equality if the learned PMF matches the true PMF exactly, $p_{\hat{\bm{y}}}(\hat{\bm{y}})=P(\hat{\bm{y}})$. Although the authors in Wagner and Ballé (2021) do not discuss the effect of the entropy model, their results on the Sawbridge imply that the learned PMF indeed can match the true (optimal) PMF exactly, resulting in no sub-optimality due to the choice of entropy model. However, this cannot be always expected, and in the case of lattice quantization, we will show that the choice of entropy model does play a role in whether optimality can be achieved.
Complexity of Entropy Estimation.
One may ask why the true entropy is not directly used. Similar to how exponential complexity is what prevents one from directly using ECVQ on general sources, it is also what prevents direct use of the true entropy $H(\hat{\bm{y}})=\operatorname{\mathbb{E}}_{\bm{x}}[-\log P(\hat{\bm{y}})]=\operatorname{\mathbb{E}}_{\hat{\bm{y}}}[-\log P(\hat{\bm{y}})]$ as a rate estimate. As mentioned previously, if operating at rate of $R$ bits per dimension, at dimension $n$, there will be on the order of $2^{nR}$ centers in the latent space. A naive plug-in estimate of the entropy will hence require at least $2^{nR}$ samples. Even the best known entropy estimators (Wu and Yang, 2016) will require $O\left(\frac{k}{\log k}\right)$ samples for an alphabet of size $k$, which is not enough of an improvement for our purposes since $k$ is on the order of $2^{nR}$. Interestingly, cross-entropy upper bounds have no such complexity issues and provide accurate estimates at large rates without requiring exponential samples (McAllester and Stratos, 2020). Hence, finding good entropy models that are able to match the true entropy closely is highly important.
Modeling General Densities.
In NTC, $p_{\bm{y}}$ is assumed factorized across its components, i.e., $p_{\bm{y}}(\bm{y})=\prod_{i=1}^{n}p_{\bm{y}_{i}}(\bm{y}_{i})$, following Ballé et al. (2021). This choice reduces modeling the density to modeling a several scalar densities, parameterized by the CDF, which is easy to integrate over the square lattice. Since we are now using a more general Monte-Carlo-based integration scheme for providing the likelihoods of $\hat{\bm{y}}$, we are no longer tied to having to parameterize $p_{\bm{y}_{i}}$ by its CDF. Instead, we propose to parameterize the PDF $p_{\bm{y}}$ in a fully multivariate fashion using a normalizing flow (Kobyzev et al., 2020). In particular, let $p_{\bm{z}}$ be $\mathcal{N}(\bm{0},\bm{I})$. Then, we model $\bm{y}=\phi(\bm{z})$ as the output of some invertible neural network $\phi$. The density of $\bm{y}$ is then
$p_{\bm{y}}(\bm{y})=p_{\bm{z}}(\phi^{-1}(\bm{y}))|\det\mathbf{J}\phi^{-1}(\bm{y})|.$ (15)
Hence, we use (15) as our flow-based entropy model. Since (15) estimates the likelihoods directly, it is more efficient to perform the Monte-Carlo integration via (15) than the factorized model, since the factorized model requires an additional differentiation of the CDF to provide likelihoods.
4.3 Vector i.i.d. Sequences
LTC can be directly applied to any vector source $\bm{x}\in\mathbb{R}^{d}$. In this case, LTC operates as a one-shot code for compressing $\bm{x}$. When $\bm{x}$ consist of i.i.d. elements $S\sim P_{S}$ on its entries, the optimal one-shot compression of $\bm{x}$ approaches the rate-distortion limit of $S$. As mentioned in the introduction, when $\bm{x}$ is a general (correlated) vector source with a low-to-moderate rate regime of interest, we can also expect coding gains when multiple i.i.d. copies $\bm{x}_{1},\dots,\bm{x}_{n}\sim p_{\bm{x}}$ are compressed together, whose optimal performance should approach $R_{\bm{x}}(D)$ when $n$ grows large. Note that we we retain $n$ as the block-length parameter, but use $d$ as the dimension of each realization of $\bm{x}$.
To realize such a compressor, we propose Block-LTC (BLTC), which applies LTC across the i.i.d. latent variables of $\bm{x}$, shown in Fig. 5.

### v2

@@ 4.1 From Scalar to Lattice Quantization
As a result, we compute $p_{\hat{\bm{y}}}(\hat{\bm{y}})$ via Monte-Carlo estimate,
@@
where $p_{\bm{u}}$ is the uniform density over $V(\hat{\bm{y}})$, $\mathrm{Vol}(V(\hat{\bm{y}}))=1$ since we assumed unit-volume lattices, and the last equality holds because for a lattice quantizer $Q_{\Lambda}(\bm{x}+\bm{v})=Q_{\Lambda}(\bm{x})+\bm{v}$, $\forall\bm{v}\in\Lambda$. From Conway & Sloane (1984), we can sample $\bm{u}^{\prime}\sim\mathrm{Unif}(V(\bm{0}))$ by computing $\bm{u}^{\prime}=\tilde{\bm{u}}-Q_{\Lambda}(\tilde{\bm{u}})$, where $\tilde{\bm{u}}=\bm{s}\bm{G}$, and $\bm{s}\sim\mathrm{Unif}([0,1])$. Thus, (5) can be estimated via a sample mean. With the ability to reliably estimate $p_{\hat{\bm{y}}}(\hat{\bm{y}})$, we can also use STE for backpropagation of the rate term $\operatorname{\mathbb{E}}_{\bm{x}}[-\log p_{\hat{\bm{y}}}(\hat{\bm{y}})]$ by using (4). Prior works have achieved benefits from using the dithering approach (Ballé et al., 2021). In this case, a random dither $\bm{u}^{\prime}\sim\mathrm{Unif}(V(0))$ is applied prior to quantization, with the resulting rate $\operatorname{\mathbb{E}}_{\bm{x},\bm{u}^{\prime}}[-\log p_{\hat{\bm{y}}}(Q_{\Lambda}(\bm{y}-\bm{u}^{\prime}))]$. This rate is equivalent in the integer lattice to $\operatorname{\mathbb{E}}_{\bm{x},\bm{u}^{\prime}}[-\log p_{\bm{y}+\bm{u}^{\prime}}(\bm{y}+\bm{u}^{\prime})]$ (Ballé et al., 2021). The density $p_{\bm{y}+\bm{u}^{\prime}}$ is a convolved density with likelihoods given by
@@
which can be estimated via Monte-Carlo. To apply this to LQ during training, rather than provide the STE of the quantized output $\hat{\bm{y}}$ to the Monte-Carlo likelihood estimate, we provide the noise-added continuous latent $\bm{y}+\bm{u}^{\prime}$, where $\bm{u}^{\prime}$ is uniformly distributed over the Voronoi region. In our experiments, we observed that STE was less stable than the dithering approach, but had better performance at low rates, as will be shown in Sec. 5.
We note that a concurrent work (Kudo et al., 2023) utilizes a similar approach to compute the rate term for the use of general lattices in image compression. Their goal was to improve coding performance by removing correlation among latent features, whereas we seek to understand the fundamental nature of quantization in nonlinear transform coding. Comparing to other related work such as LVQAC (Zhang & Wu, 2023), we show later how their use of $D_{n}^{*}$ lattice results in fundamental suboptimality.
Entropy Models. The rate term, $\operatorname{\mathbb{E}}_{\bm{x}}[-\log p_{\hat{\bm{y}}}(\hat{\bm{y}})]$, which is the cross-entropy between the true PMF on the lattice vectors $P(\hat{\bm{y}}):=\Pr_{\bm{x}}(Q_{\Lambda}(g_{a}(\bm{x})=\hat{\bm{y}})$ and $p_{\hat{\bm{y}}}(\hat{\bm{y}})$, is an upper bound on the true entropy of $P(\hat{\bm{y}})$, $\operatorname{\mathbb{E}}_{\bm{x}}[-\log P(\hat{\bm{y}})]$, with equality if the learned PMF matches the true PMF exactly, $p_{\hat{\bm{y}}}(\hat{\bm{y}})=P(\hat{\bm{y}})$. Using the cross-entropy as an estimate of the true entropy is critical due to the limitations of entropy estimation. Whereas direct entropy estimation requires a number of samples exponential in the rate, cross-entropy upper bounds are able to provide accurate estimates at large rates without requiring exponential samples (McAllester & Stratos, 2020). We cannot always expect the bound to be tight, and we will show that the choice of entropy model does play a role in whether optimality can be achieved in Sec. 5.3. Since we are now using a more general Monte-Carlo integration for the likelihoods of $\hat{\bm{y}}$, we are no longer tied to a factorized density and instead propose to parameterize the PDF $p_{\bm{y}}$ in a multivariate fashion using a normalizing flow (Kobyzev et al., 2020).
4.2 Toward Lower Complexity
However, as the dimension of the lattice increases, entropy coding becomes challenging due to the large alphabet size. Large alphabet entropy coding is indeed an active area of research. Another approach that circumvents the need for entropy coding altogether is by quantization using nested lattices. To implement a fixed-rate code in the latent space using lattices, we use a nested lattice in the latent space. The idea is to use the Voronoi region of a coarse lattice $\Lambda_{c}$ to define a bounded region of a fine lattice $\Lambda_{f}$. The fixed number of codebook vectors of $\Lambda_{f}$ contained in the said region are the ones we wish to use to encode the source via index coding (and thus a fixed rate). This happens when the majority of the latent mass is placed in the Voronoi region of the coarse lattice. Efficient encoding/decoding is enabled by transmitting the index of the coset, yielding a complexity of $O(n)$ for an alphabet of size $O(2^{n})$, compared to entropy coding which will be $O(2^{n})$. To quantize the latent $\bm{y}$, we compute $\hat{\bm{y}}=\bm{y}_{f}-Q_{\Lambda_{c}}(\bm{y}_{f})$, where $\bm{y}_{f}=Q_{\Lambda_{f}}(\bm{y})$. This operation selects the coset leader corresponding to $\bm{y}_{f}$ inside the Voronoi region of $\Lambda_{c}$. The rate is thus given by $\log_{2}|\Lambda_{1}/\Lambda_{2}|=\Gamma^{d_{y}}$, where $|\Lambda_{1}/\Lambda_{2}|$ is the number of relative cosets, $d_{y}$ is the dimension of $\bm{y}$, and $\Gamma$ is the nesting ratio (Zamir et al., 2014).
The majority of the latent mass is placed in the Voronoi region of $\Lambda_{c}$ when $Q_{\Lambda_{c}}(\bm{y}_{f})=\bm{0}$ occurs with high probability. Otherwise, we have an error event :=$\mathcal{E}:=\{Q_{\Lambda_{c}}(\bm{y}_{f})\neq\bm{0}\}$. Let $P_{e}=\Pr(\mathcal{E})$. The total distortion can be broken up into two terms (Zamir et al., 2014, Ch. 10),
$\operatorname{\mathbb{E}}_{\bm{x}}[\operatorname{\mathsf{d}}(\bm{x},\hat{\bm{x}})]=(1-P_{e})\cdot\operatorname{\mathbb{E}}_{\bm{x}}\left[\operatorname{\mathsf{d}}(\bm{x},\hat{\bm{x}})|\mathcal{E}^{\complement}\right]+P_{e}\cdot\operatorname{\mathbb{E}}_{\bm{x}}\left[\operatorname{\mathsf{d}}(\bm{x},\hat{\bm{x}})|\mathcal{E}\right].$ (7)
In order to train LTC with latent NLQ, we need to be able to differentiate through these terms. We use a normalizing flow density $p_{\bm{y}}$ to estimate the $P_{e}$ with Monte-Carlo integration,
@@ 4.2 Toward Lower Complexity
where $V_{c}(\bm{0})$ is the Voronoi region of $\Lambda_{c}$. To enforce $p_{\bm{y}}$ to model the true distribution of $\bm{y}=g_{a}(\bm{x})$, we maximize its log likelihood $\operatorname{\mathbb{E}}_{\bm{x}}[\log p_{\bm{y}}(g_{a}(\bm{x}))]$. Thus, our complete loss is given by
$\min_{g_{a},g_{s},p_{\bm{y}}}(1-\hat{P}_{e})\cdot\operatorname{\mathbb{E}}_{\bm{x}}\left[\operatorname{\mathsf{d}}(\bm{x},\hat{\bm{x}})|\mathcal{E}^{\complement}\right]+\hat{P}_{e}\cdot\operatorname{\mathbb{E}}_{\bm{x}}\left[\operatorname{\mathsf{d}}(\bm{x},\hat{\bm{x}})|\mathcal{E}\right]+\operatorname{\mathbb{E}}_{\bm{x}}[-\log p_{\bm{y}}(g_{a}(\bm{x}))].$ (9)
In our experiments, we found that detaching the gradient from the $\operatorname{\mathbb{E}}_{\bm{x}}\left[\operatorname{\mathsf{d}}(\bm{x},\hat{\bm{x}})|\mathcal{E}\right]$ term helped to significantly improve training stability. For the lattices used, we used self-similar nested lattices: $G_{c}=\Gamma\cdot G_{f}$, where $G_{c}$ and $G_{f}$ are the coarse and fine lattices’ generator matrices respectively, and $\Gamma$ is the integer-valued nesting ratio, used to vary the rate of the compressor.
4.3 From One-Shot to Block Coding